# Reference map technique for Lagrangian exploration of coherent structures

### v1

@@ 3 Results and discussions
Figure 4 compares the backward-time FTLE ($\Lambda^{t_{0}}_{t}(\textbf{x})$) obtained with the reference map to that obtained with a direct particle-based approach. The latter is obtained from a 2D open-source MATLAB program, LCS Tool (Onu et al., 2015). For LCS Tool, the Lagrangian tracer resolution is identical to the grid resolution for the reference map. It is readily confirmed from the left and middle panels of figure 4 that the FTLE fields produced by the particle-based method and the reference map approach are indistinguishable, which serves as validation of the current approach. As ridges (local maximizing lines/surfaces) of the backward-time FTLE are candidates for attracting material lines or attracting LCS (Haller & Yuan, 2000), we further attempt to locate the FTLE ridges through the zero-gradient contours of the FTLE that satisfy a second-derivative condition in the $x$ direction.
@@
Figure 2: 2D steady Taylor-Green Vortex: Time evolution of contour of $x_{2}$-component of the reference map ($\xi_{2}(x,y,t)/(2\pi)$) from $t=0$ (top left) to $t=22$ (bottom right) (time interval =2).
@@
Figure 4: 2D steady Taylor-Green Vortex: Comparison of color contours of the largest backward-time FTLE ($\Lambda^{t_{0}}_{t}$) obtained from $t=10$ to $t_{0}=0$. Left, from LCS Tool (particle-based approach). Middle, from the reference map (present study). Right, Same as Middle, but the solid lines are the FTLE ridges as identified from Eq. (14). Grid/particle resolutions are the same in all cases (512x512).
@@ 3.2 Turbulent Channel flow
The DNS of turbulent channel flow at $Re_{\tau}=180$ was performed, where $Re_{\tau}$ is the friction Reynolds number based on the channel half-height ($\delta$), friction velocity ($u_{\tau}$), and kinematic viscosity ($\nu$). A minimal channel with dimensions $x/\delta\in[0,2\pi]$ in the streamwise direction, $y/\delta\in[-1,1]$ in the wall-normal direction, and $z/\delta\in[0,\pi]$ in the spanwise direction was used, where $x$, $y$, and $z$ represent the streamwise, wall-normal, and spanwise directions, respectively. Periodicity was imposed in the streamwise and spanwises directions. A uniform grid was used in the streamwise and the spanwise directions, whereas a stretched grid was employed in the wall-normal direction, with the number of grid points in each direction given by $(N_{x},N_{y},N_{z})=(200,260,160)$. The stretching rate for the wall-normal spacing is given by the geometric progression $\Delta y_{i}=r^{i-1}\Delta y_{1}$, $i=1,2,...,N_{y}/2$, where $r=1.035$ and $\Delta y_{1}/\delta=4.04\times 10^{-4}$. The resulting grid spacing in wall units are $(\Delta x^{+},\Delta y_{min}^{+},\Delta y_{max}^{+},\Delta z^{+})\approx(6,0.07,6,4)$. After the simulation reached statistical stationarity, statistics were collected over a nondimensional time $T^{+}=tu^{2}_{\tau}/\nu\approx 9500$ ($tU_{b}/\delta\approx 822$). Profiles of the mean velocity and the Reynolds stresses from the present DNS were validated against the channel flow DNS of Kim et al. (1987). Computation of the FTLE field started at $t=t_{0}$ once the flow was fully developed, where The reference map variable was set to be the coordinates of cell centers.
@@
Figure 5: Comparison of backward-time FTLE with $Q$-criteria structures at $t^{+}=12+t_{0}^{+}$. In the sub-figures where available, the iso-surfaces colored by instantaneous streamwise velocity are those of $Q^{+}=350$, which are shown along with $(a)$ iso-surfaces of $\Lambda^{t_{0}}_{t}=20$, $(b)$ iso-surfaces of $\Lambda^{t_{0}}_{t}\leq 15$, $(c)$ backward-time FTLE in a constant-streamwise cut at $x/\delta=6$, $(d)$ backward-time FTLE in a constant-spanwise cut at $z/\delta=1.5$; $(e)$ $(g)$ show $\Lambda^{t_{0}}_{t}$ in the same constant-streamwise cut as $(c)$; $(f)$ shows $\Lambda^{t_{0}}_{t}$ in the same constant-spanwise cut as $(d)$. In $(e)$ $(f)$, the red boxes identify structures to be compared with those from Green et al. (2007) shown in the insets. In $(g)$, the colored arrows indicate the direction of in-plane instantaneous velocity vector (colored by $Q$-criterion).
@@
Using the procedure described in section 2.1, the backward-time FTLE field is evaluated at $t^{+}=12+t_{0}^{+}$. At this time instance, figure 5 compares the backward-time FTLE with Q-criterion by visualizing a couple of hairpin structures at the wall. As the potential attracting LCSs are identified with the ridges (local maxima) of the backward-time FTLE, they are commonly visualized through two-dimesional planes rather than three-dimensional iso-surfaces (Green et al., 2007). Nonetheless, we first look at iso-surfaces of the backward-time FTLE to get a global picture of the FTLE field. Figure 5$(a)$ shows that the iso-surfaces of $\Lambda^{t}_{t_{0}}=20$ approximately coincide with the hairpin vortices identified by iso-surface of Q-criterion. However, a single FTLE iso-surface does not faithfully depict the boundaries of Lagrangian structures, and it is more appropriate to visualize a range of constant-FTLE surfaces (or volume) above a prescribed high value as in Green et al. (2007). Conversely, the negative volume etched by plotting constant-FTLE surfaces below a prescribed FTLE limit can be visualized and compared with the Q-criterion structures. Figure 5$(b)$ utilizes this visualization technique by plotting the volume of the region with $\Lambda^{t_{0}}_{t}\leq 15$. It is observed (from multiple view angles although not shown all here) that the hairpin vortices from Q-criterion fill up the negative region created by this volume, showing that the boundaries of coherent structures can indeed be delineated using FTLE. In figure 5$(c)$ and $(d)$, FTLE is plotted in a streamwise and spanwise cut-plane, respectively, to illustrate the two-dimensional skeleton of LCS. It is clear that the regions of high FTLE closely correlate with the vortical structures given by the iso-surfaces of $Q$-criteria. The “mushroom” structure in figure 5$(e)$, and the roll-up structure in figure 5$(f)$ are indeed reminiscent of LCS computed from a particle approach in Green et al. (2007) for the same flow, as shown in the insets of these figures. Further evidence of the correlation between FTLE and $Q$-criteria structures is provided by figure 5$(g)$, where the vortices indicated by the swirling velocity vectors (also corresponding to the regions of high $Q$) tend to be bounded by regions of locally high $\Lambda^{t}_{t_{0}}$, consistent with the finding of Green et al. (2007). Specifically, for the “mushroom” structure in figure 5$(g)$, the head of the mushroom closely wraps around the pair of counter-rotating vortices corresponding to the legs of the hairpin structure. Closer to the wall (to the bottom left and right of the mushroom), we observe more counter-rotating vortices likely related to the formation of low-speed streaks. This will be further evidenced through the visualization of material surfaces in figure 8.
@@
A rather rudimentary but insightful way to to visualize fluid element deformation at a macroscopic level is through the evolution of initially planar material surfaces (Yeung, 2002), which is provided directly by the reference map $\bm{\xi}$. Specifically, we track how the tracers that form a constant-streamwise and a constant-wall-normal plane at $t=0$ evolve in response to the background turbulent flow. Note that this is essentially the 3D counterpart of material-line tracking in figure 3. Figure 6 shows the time evolution of the plane identified by $\xi_{1}(x,y,z,t=0)=1.5\delta$ from $t^{+}=0$ to $t^{+}=4$. It is interesting to note that the structures visible from the instantaneous velocity at $t^{+}=0$ leave a mark on the material surface as it evolves in time. Specifically, a mushroom structure indicative of a hairpin vortex is observed toward the top right of the plane. The regions of lower velocity within this structure lag behind the higher-velocity fluid particles outside it, leaving an imprint of the background turbulent flow structures on the material surface at $t^{+}=4$. In the right panel of figure 6, the backward-time FTLE is plotted on the deformed $\xi_{1}(x,y,z,t)$ surface showing that the imprinted structures correspond to regions of high FTLE.
@@
To get a more holistic and consolidated picture of the flow structures identified individually from the $x$- and $y$-plane tracking, in figure 8 we superimpose at $t^{+}=8$ the constant $\xi_{1}$ and constant $\xi_{2}$ surfaces initially at $x/\delta=2.4$ and $y/\delta=-0.85$, respectively. Furthermore, to show the relation of these structures to those from $Q$-criterion, the velocity vectors colored by $Q$-criterion are plotted onto the deformed $\xi_{1}$ surface. We focus on one particular instance of a structure shown in the box. Here, the $\xi_{2}$ surface indicates the presence of an elongated low-speed streak through elevation in the material surface. Consistent with the well-studied mechanism for the low-speed streaks, this streak lies exactly between a pair of counter-rotating vortices depicted by the velocity vectors and the regions of high $Q$. The velocity vectors clearly show the ejection of fluid away from the wall between the two vortices. Furthermore, the lagging region of $\xi_{1}$ surface (described above in relation to figure 6) also coincides with the low-speed streak and is located between the counter-rotating vortex pair. The $\xi_{1}$ material surface exhibits twisting which is consistent with the direction of each of the counter-rotating vortices.
@@ 4 Discussion and Conclusion
The information implicitly embedded in the reference map regarding fluid particle trajectories provides a means to calculate measures of Lagrangian fluid element deformation solely from Eulerian data. These measures encompass displacement, deformation gradient tensor, and (left/right) Cauchy-Green tensor. They can be further leveraged to compute the backward-time Finite-Time Lyapunov Exponent (FTLE) field, with its ridges identified as candidates for the attracting LCS. We demonstrated the accuracy of FTLE calculations based on the reference map in comparison to the standard particle-based approach using a simple two-dimensional flow (2D Taylor-Green vortex). LCS detection utilizing the reference map was then applied to turbulent channel flow at $Re_{\tau}=180$. The computed LCSs exhibited consistency with structures identified through the $Q$ criterion, where the FTLE ridges closely approximated the boundaries of the $Q$ structures, as illustrated in Green et al. (2007). A notable drawback of the current approach, in contrast to widely used Eulerian vortex identification schemes, is the high cost associated with incorporating an additional three advection equations for the reference map. While Eulerian schemes reveal coherent structures based on the instantaneous velocity field, LCS based on the concept of finite-time dynamical systems necessitates fluid particle trajectories over a specified time interval. Given that dominant LCSs tend to converge over relatively short time scales (Huang et al., 2022; Green et al., 2007), this increased computational cost may be confined to a small fraction of the total simulation cost by integrating the reference map equation only during the time it is needed.

### v2

@@ 4 Results and discussions
Figure 6 compares the backward-time FTLE ($\Lambda^{t_{0}}_{t}(\textbf{x})$) obtained with the reference map to that obtained with a direct particle-based approach. For LCS Tool, the main particle grid for Lagrangian particle tracking is identical to the grid for the reference map (a $512\times 512$ uniformly-spaced grid). However, LCS Tool utilizes an auxiliary particle grid (4 points surrounding each main grid point) with the spacing of 1–10% of the main grid spacing to approximate the deformation gradient accurately with finite difference (Onu et al., 2015). It is readily confirmed from the left and middle panels of figure 6 that the FTLE fields produced by the particle-based method and the reference map approach are indistinguishable, which serves as validation of the current approach. As ridges (local maximizing lines/surfaces) of the backward-time FTLE are candidates for attracting material lines or attracting LCS (Haller & Yuan, 2000), we further attempt to locate the FTLE ridges through the zero-gradient contours of the FTLE that satisfy a second-derivative condition in the $x$ direction
@@
Additionally, we perform a convergence study of the reference map based approach under grid refinement, and take the backward-time FTLE field computed on a fine grid $(1024\times 1024)$ using the reference map approach as a reference solution. Specifically, we consider the following uniform grids $64\times 64$, $128\times 128$, $256\times 256$, $512\times 512$, and compute relative $L^{2}$ errors against the fine-grid result. Figure 7(a) plots these errors as a function of uniform grid spacing, demonstrating the reference map backward-time FTLE field converges at a rate of approximately 0.7. Next, we assess the difference between the backward-time FTLE fields computed on the same grid using the reference map approach and LCS Tool. The difference, albeit small, is concentrated most in the boundaries between the neighboring spiral vortices where rapid variations in the strain rate and FTLE are observed, as seen in the normalized absolute difference plot in Figure 7(b). As noted earlier, the use of the auxiliary particle grid results in extra accuracy in LCS Tool, which is absent in the current reference map approach. Nevertheless, we observed the difference decreases under grid refinement.
Figure 4: 2D steady Taylor-Green Vortex: Time evolution of contour of $x_{2}$-component of the reference map ($\xi_{2}(x,y,t)/(2\pi)$) on the periodic domain $[0,2\pi]^{2}$ from $t=0$ to $t=10$ in the top row and $t=12$ to $t=22$ in the bottom row (time interval =2).
@@
Figure 6: 2D steady Taylor-Green Vortex: Comparison of color contours of the largest backward-time FTLE ($\Lambda^{t_{0}}_{t}$) obtained from $t=10$ to $t_{0}=0$. Left, from LCS Tool (particle-based approach). Middle, from the reference map (present study). Right, Same as Middle, but the solid lines are the FTLE ridges as identified from Eq. (17). Grid/particle resolutions are the same in all cases ( $512\times 512$).
Figure 7: 2D steady Taylor-Green Vortex: (a) Relative $L^{2}$ error between a between a fine grid ($1024\times 1024$) reference backward-time FTLE field computed using the RMT approach and those on a sequence of successively refined grids, where we obtain a convergence rate of approximately $h^{0.7}$. (b) Normalized absolute difference between the backward-time FTLE fields computed using the particle-based (LCS Tool) and the reference-map approaches on a 512$\times$512 grid.
4.2 2D Lid-driven cavity
The next case we consider for the validation of reference-map-based FTLE calculation is an incompressible two-dimensional steady lid-driven cavity at a Reynolds number of 100. The geometry is a square cavity with unit edge length, consisting of three rigid walls with no-slip conditions and the top wall moving along positive $x$ with a tangential unit velocity, inducing the flow inside the cavity. At $t=0$ where the flow has reached the steady state, the reference map variables are initialized with the coordinates of grid collocation points. Similar to the 2D TG vortex, the velocity component normal to all the walls is 0, simplifying the boundary closure of the reference map as discussed in Sec. 3.
In Fig. 8, the steady velocity profiles at the mid-sectional lines $x=0.5$ and $y=0.5$ from a simulation using 512$\times$512 grid are validated against the profiles from Ghia et al. (1982). The profiles of both horizontal and vertical velocity show good agreement with the reference data. Figure 9 shows the comparison of the backward-time FTLE ($\Lambda^{t_{0}}_{t}(\textbf{x})$) from the present reference-map approach and that from a direct particle-based approach using the LCS Tool described in the previous section, both computed on a $512\times 512$ uniformly-spaced grid. For the Lagrangian tracking in the LCS Tool, a fine auxiliary grid is used with a grid spacing equal to 1% of the main grid spacing. A visual inspection of the FTLE fields produced by the LCS Tool and the reference map approach in figure 9, shows a reasonable overall agreement between the two fields, except for some localized regions of high $\Lambda^{t_{0}}_{t}(\textbf{x})$ toward the bottom right of the right panel. These localized regions appear to be originating from the singularity of the boundary condition at the top-right corner (Bouffanais et al., 2007; Sousa et al., 2016; Kuhlmann & Romanò, 2019). This effect is not observed in the velocity field but only in the FTLE field. Note that the scalar advection equation (which the reference map obeys) is more prone to the dispersion error due to such discontinuous data. However, we find that the local nature of these artifacts prevent them from significantly polluting most of the FTLE field, even after a long time integration of 11 time unit.
Figure 8: Validation of velocity profiles in 2D lid-driven cavity at Re=100: horizontal velocity at $x=0.5$ (left) and vertical velocity at $y=0.5$ (right). Squares, reference data from Ghia et al. (1982); solid lines, present simulation on a 512$\times$512 grid.
Figure 9: 2D lid-driven cavity at Re=100: Comparison of color contours of the largest backward-time FTLE ($\Lambda^{t_{0}}_{t}$) obtained from $t=11.24$ to $t_{0}=0$. Left, from LCS Tool (particle-based approach). Right, from the reference map (present study).
@@ 4.3 Turbulent Channel flow
The DNS of turbulent channel flow at $Re_{\tau}=180$ was performed, where $Re_{\tau}$ is the friction Reynolds number based on the channel half-height ($\delta$), friction velocity ($u_{\tau}$), and kinematic viscosity ($\nu$). A minimal channel with dimensions $x/\delta\in[0,2\pi]$ in the streamwise direction, $y/\delta\in[-1,1]$ in the wall-normal direction, and $z/\delta\in[0,\pi]$ in the spanwise direction was used, where $x$, $y$, and $z$ represent the streamwise, wall-normal, and spanwise directions, respectively. Periodicity was imposed in the streamwise and spanwise directions. A uniform grid was used in the streamwise and spanwise directions, whereas a stretched grid was employed in the wall-normal direction, with the number of grid points in each direction given by $(N_{x},N_{y},N_{z})=(200,260,160)$. The wall-normal spacing is given by the geometric progression $\Delta y_{i}=r^{i-1}\Delta y_{1}$, $i=1,2,...,N_{y}/2$, where $r=1.035$ and $\Delta y_{1}/\delta=4.04\times 10^{-4}$. The resulting grid spacings in wall units are $(\Delta x^{+},\Delta y_{min}^{+},\Delta y_{max}^{+},\Delta z^{+})\approx(6,0.07,6,4)$. Note that the superscript “$+$” denotes quantities non-dimensionalized by viscous wall units, where the non-dimensionalized distance $l^{+}$ and time $t^{+}$ are given by, $l^{+}=lu_{\tau}/\nu$ and $t^{+}=tu^{2}_{\tau}/\nu$, respectively. After the simulation reached statistical stationarity (which is verified through the convergence of integrated wall forces and bulk velocity), statistics were collected over a nondimensional time period $T^{+}=tu^{2}_{\tau}/\nu\approx 9500$ ($tU_{b}/\delta\approx 822$ , where $U_{b}$ is the bulk velocity ). Profiles of the mean velocity and the Reynolds stresses from the present DNS were validated against the channel flow DNS of Kim et al. (1987). After the flow was fully developed, computation of the FTLE field was started at an arbitrary time $t=t_{0}$, where the reference map variable was set to be the coordinates of cell centers.
@@
Figure 10: Comparison of the backward-time FTLE with the $Q$-criterion structures at $t^{+}=12+t_{0}^{+}$. In the sub-figures where available, the iso-surfaces colored by instantaneous streamwise velocity are those of $Q^{+}=350$, which are shown along with $(a)$ iso-surfaces of $\Lambda^{t_{0}}_{t}\delta/u_{\tau}=20$, $(b)$ iso-surfaces of $\Lambda^{t_{0}}_{t}\delta/u_{\tau}\leq 15$, $(c)$ backward-time FTLE in a constant-streamwise cut at $x/\delta=6$, $(d)$ backward-time FTLE in a constant-spanwise cut at $z/\delta=1.5$; $(e)$ $(g)$ show $\Lambda^{t_{0}}_{t}$ in the same constant-streamwise cut as $(c)$; $(f)$ shows $\Lambda^{t_{0}}_{t}$ in the same constant-spanwise cut as $(d)$. In $(e)$ $(f)$, the red boxes identify structures to be compared with those from Green et al. (2007) shown in the insets. Note that Green et al. (2007) used a particle-tracking (auxiliary) grid for the FTLE calculation that was six-times finer in all directions than the flow grid for DNS. In $(g)$, the colored arrows indicate the direction of in-plane instantaneous velocity vector (colored by the $Q$-criterion).
@@
Using the procedure described in section 2.1, the backward-time FTLE field is evaluated at $t^{+}=12+t_{0}^{+}$. At this time instance, figure 10 ($a-d$) compare the backward-time FTLE with the Q-criterion by visualizing a couple of hairpin structures at the wall. As the potential attracting LCSs are identified with the ridges (local maxima) of the backward-time FTLE, they are commonly visualized through two-dimesional planes rather than three-dimensional iso-surfaces (Green et al., 2007). Nonetheless, we first look at iso-surfaces of the backward-time FTLE to get a global picture of the FTLE field. Figure 10$(a)$ shows that the iso-surfaces of $\Lambda^{t_{0}}_{t}\delta/u_{\tau}=20$ approximately coincide with the hairpin vortices identified by iso-surface of the Q-criterion. However, a single FTLE iso-surface does not faithfully depict the boundaries of Lagrangian structures, and it is more appropriate to visualize a range of constant-FTLE surfaces (or volume) above a prescribed high value as in Green et al. (2007). Conversely, the negative volume etched by plotting constant-FTLE surfaces below a prescribed FTLE limit can be visualized and compared with the Q-criterion structures. Figure 10$(b)$ utilizes this visualization technique by plotting the volume of the region with $\Lambda^{t_{0}}_{t}\delta/u_{\tau}\leq 15$. It is observed (from multiple view angles although not all shown here) that the hairpin vortices from the Q-criterion fill up the negative region created by this volume, showing that the boundaries of coherent structures can indeed be delineated using FTLE. In figure 10$(c)$ and $(d)$, FTLE is plotted in a streamwise and spanwise cut-plane, respectively, to illustrate the two-dimensional skeleton of LCS. It is clear that the regions of high FTLE closely correlate with the vortical structures given by the iso-surfaces of the $Q$-criterion . The “mushroom” structure in figure 10$(e)$, and the roll-up structure in figure 10$(f)$ are indeed reminiscent of LCS computed from a particle approach in Green et al. (2007) for the same flow, as shown in the insets of these figures. Further evidence of the correlation between the FTLE and $Q$-criterion structures is provided by figure 10$(g)$, where the vortices indicated by the swirling velocity vectors (also corresponding to the regions of high $Q$) tend to be bounded by regions of locally high $\Lambda^{t}_{t_{0}}$, consistent with the finding of Green et al. (2007). Specifically, for the “mushroom” structure in figure 10$(g)$, the head of the mushroom closely wraps around the pair of counter-rotating vortices corresponding to the legs of the hairpin structure. Closer to the wall (to the bottom left and right of the mushroom), we observe more counter-rotating vortices likely related to the formation of low-speed streaks. This will be further evidenced through the visualization of material surfaces in figure 13.
@@
A rather rudimentary but insightful way to visualize fluid element deformation at a macroscopic level is through the evolution of initially planar material surfaces (Yeung, 2002), which is provided directly by the reference map $\bm{\xi}$. Specifically, we track how the tracers that form a constant-streamwise and a constant-wall-normal plane at $t=0$ evolve in response to the background turbulent flow. Note that this is essentially the 3D counterpart of material-line tracking in figure 5. Figure 11 shows the time evolution of the plane identified by $\xi_{1}(x,y,z,t=0)=1.5\delta$ from $t^{+}=0$ to $t^{+}=4$. It is interesting to note that the structures visible from the instantaneous velocity at $t^{+}=0$ leave a mark on the material surface as it evolves in time. Specifically, a mushroom structure indicative of a hairpin vortex is observed toward the top right of the plane. The regions of lower velocity within this structure lag behind the higher-velocity fluid particles outside it, leaving an imprint of the background turbulent flow structures on the material surface at $t^{+}=4$. In the right panel of figure 11, the backward-time FTLE is plotted on the deformed $\xi_{1}(x,y,z,t)$ surface showing that the imprinted structures correspond to regions of high FTLE.
@@
To get a more holistic and consolidated picture of the flow structures identified individually from the $x$- and $y$-plane tracking, in figure 13 we superimpose at $t^{+}=8$ the constant $\xi_{1}$ and constant $\xi_{2}$ surfaces initially at $x/\delta=2.4$ and $y/\delta=-0.85$, respectively. Furthermore, to show the relation of these structures to those from the $Q$-criterion, the velocity vectors colored by the $Q$-criterion are plotted onto the deformed $\xi_{1}$ surface. We focus on one particular instance of a structure shown in the box. Here, the $\xi_{2}$ surface indicates the presence of an elongated low-speed streak through elevation in the material surface. Consistent with the well-studied mechanism for the low-speed streaks, this streak lies exactly between a pair of counter-rotating vortices depicted by the velocity vectors and the regions of high $Q$. The velocity vectors clearly show the ejection of fluid away from the wall between the two vortices. Furthermore, the lagging region of $\xi_{1}$ surface (described above in relation to figure 11) also coincides with the low-speed streak and is located between the counter-rotating vortex pair. The $\xi_{1}$ material surface exhibits twisting which is consistent with the direction of each of the counter-rotating vortices.
@@ 5 Discussion and Conclusion
The information implicitly embedded in the reference map regarding fluid particle trajectories provides a means to calculate measures of Lagrangian fluid element deformation solely from Eulerian data. These measures encompass displacement, deformation gradient tensor, and (left/right) Cauchy-Green tensor. They can be further leveraged to compute the backward-time Finite-Time Lyapunov Exponent (FTLE) field, with its ridges identified as candidates for the attracting LCS. We demonstrated the accuracy of FTLE calculations based on the reference map in comparison to the standard particle-based approach using a simple two-dimensional flow (2D Taylor-Green vortex). LCS detection utilizing the reference map was then applied to turbulent channel flow at $Re_{\tau}=180$. The computed LCSs exhibited consistency with structures identified through the $Q$-criterion, where the FTLE ridges closely approximated the boundaries of the $Q$-criterion structures, as illustrated in Green et al. (2007). A notable drawback of the current approach, in contrast to widely used Eulerian vortex identification schemes, is the high cost associated with incorporating an additional three advection equations for the reference map. While Eulerian schemes reveal coherent structures based on the instantaneous velocity field, LCS based on the concept of finite-time dynamical systems necessitates fluid particle trajectories over a specified time interval. Given that dominant LCSs tend to converge over relatively short time scales (Huang et al., 2022; Green et al., 2007), this increased computational cost may be confined to a small fraction of the total simulation cost by integrating the reference map equation only during the time it is needed.